### v1

# Ab Initio Free Energy Surfaces for Coupled Ion-Electron Transfer

Abstract
The Marcus theory of electron transfer assumes that diabatic energy gaps are sampled from a single ensemble. This assumption can break down in spatially anisotropic environments, such as electrochemical interfaces or biomolecular structures, where distinct solvent ensembles arise along a collective variable describing the anisotropy. Treating this collective variable as an additional reaction coordinate linearly independent from the Marcus reaction coordinate, we develop a formalism that enables calculation of the resulting Coupled Ion–Electron Transfer (CIET) free-energy surface directly from constrained ab initio trajectories. Applied to CO2 redox on a gold electrode, this method reveals strong coupling to the anisotropy, predicting significantly different activation barriers compared to either coordinate alone.
Marcus free energy curves relate macroscopically observed electron transfer (ET) reaction rates to a microscopic picture of solvent fluctuations [1, 2, 3]. The iconic parabolic free energy curves that describe the ET landscape arise from the fact that the energy gap between the reactant and product diabatic electronic states is mediated by the instantaneous nuclear configurations of the solvation environment, and that the probability distribution of such energy gaps is Gaussian to a good approximation; taking the logarithm of this distribution yields the parabolas [4, 5, 6]. Crucially, this framework assumes that nuclear configurations are sampled from a single, global equilibrium ensemble—a simplification that may break down in anisotropic environments such as electrochemical interfaces or biomolecular structures [7, 8, 9, 10].
In this Letter, we show that in the presence of such anisotropy, ET can become coupled to a classical collective variable (CV) that characterizes the broken symmetry. This is suggested by the fact that the nuclear ensembles, when conditioned on different values of the CV, are generally distinct. As a result, the Marcus curves associated with each value may differ in curvature or vertical shift, and the reaction landscape is more appropriately viewed as a 2D manifold of Marcus curves parametrized by the anisotropic CV.
Extensions of Marcus theory to cases in which ET is coupled to another degree of freedom have already found widespread applications. Proton-coupled electron transfer (PCET) generalizes Marcus theory by extending the reaction coordinate to include a proton degree of freedom [9, 10]. However, the proton is typically treated quantum mechanically and associated with a reaction coordinate involving vibronic excitations, precluding a straightforward generalization to the case when ET is coupled to a classical CV. More recently, coupled ion-electron transfer (CIET), proposed by Bazant and coworkers [7, 8], has emerged as a framework to explain the rates of redox reactions in commercial lithium-ion battery electrodes [11, 12]. In this theory, the Marcus ET coordinate is coupled to a classical ion transfer (IT) coordinate related to the distance from the electrode-electrolyte interface, and the resulting rate expressions fit the experimentally observed current-voltage curves better than rate expressions that do not take into account ion-electron coupling [13, 12, 7].
Despite the emergence of Marcus theory in recent decades as a physically correct way to view ET reactions, the Butler–Volmer (BV) equation, which has empirically expressed Tafel’s law for current-voltage relationships of reaction-limited currents for nearly a century, assumes diabatic free energy curves that are linear instead of quadratic [14, 15]. In CIET, this discrepancy is resolved by postulating a 2D free energy surfaces which can in general have an arbitrary structure, but in the simplest approximation is linear when cross-sectioned along the ion-transfer CV and quadratic along the Marcus ET coordinate [7]. The result is that the CIET rate expression reduces to the Marcus–Hush–Chidsey (MHC) equation [16, 17, 18] in the limit where the reorganization energy dominates, and reduces to the BV equation in the limit where the barrier along the classical IT coordinate dominates [7]. The fact that viewing an electron transfer reaction as coupled to an anisotropic CV not only explains experimental reaction rates but also unifies physical equations correct in different regimes suggests the utility of such an approach, and suggests the possibility that a related framework can be generalized to symmetry-broken CVs other than IT at intercalation cathodes, finding broader applications across condensed phase physical chemistry [19, 20]. Practical analyses using these multidimensional surfaces may also become more feasible with concurrent improvements in sampling methods and machine-learned potentials [21, 22, 23, 24, 25, 26].
We note that ab initio computation of CIET parameters has recently been pioneered by Stenlid et al. [11] on a model lithium-cobalt-oxide battery interface. While remarkably effective in qualitatively reproducing experimental kinetic data, these calculations were conducted at 0 K, leaving entropic contributions to the free energy unaddressed. Furthermore, the study stopped short of computing the full 2D landscape, making assumptions, for example, that the reorganization energy was independent of the IT coordinate, and that adiabatic corrections to the landscape were insignificant.
Here we present a formalism to compute finite-temperature 2D free-energy surfaces for ET coupled to a classical CV, expressed entirely in terms of quantities derived from constrained ab initio trajectories. We incorporate adiabatic corrections to the landscape as well, which prove to be significant [16]. We illustrate the method by applying it to CO2 redox on a gold electrode, and we find that the transition state (TS) of the 2D landscape can be significantly lower than when computed from the ET coordinate or CV alone. Although this demonstration assumes harmonic and symmetric ET, the framework allows this assumption to be readily relaxed, enabling generalization to diverse systems.
Constructing the 2D Free Energy Surface. Within the Born–Oppenheimer approximation, an ab initio trajectory constrained to a diabatic electronic state can be modeled using constrained density functional theory (cDFT), as developed by Van Voorhis and co-workers [27, 28, 29] and implemented in various software packages [30, 31, 32]. Consider a simulation of atoms with the electronic state fixed to , where denotes the nuclear configuration. Let denote the nuclear trajectory under this constraint. At steady state, the probability of finding the system in configuration is given by the time-independent distribution . Letting denote the initial and final (i.e. oxidized and reduced) diabatic states of the electron transfer, the ET coordinate of is defined by the energy gap [4, 33, 34]
where is the Hamiltonian. Now suppose is a CV that characterizes the anisotropy. Then the histogram is obtained from the trajectory by integrating over the appropriate region of the phase space
The the probability distribution is then given by , and the 2D excess free energy 111Formally, the free energy used in this work is the excess free energy , which retains only the terms that affect the potential experienced by an individual particle, subtracting off the ”counting” effect related to concentration [7]. However, we denote this quantity by the standard free energy symbol for notational simplicity. surface is given by
where is a constant. It is convenient when applicable to define the equilibrium coordinates for each state
Then, noting that the constants and must be related such that the reaction free energy between these equilibrium states is the thermodynamic driving force , a natural choice is
so that the equilibrium reactant state has .
A practical challenge is that in the presence of high reaction barriers, the sampled histogram is often unpopulated for many regions in the 2D reaction coordinate space including, crucially, near the TS. However, this challenge can easily be overcome by employing methods common in Marcus theory (for ) and classical enhanced sampling (for ). Sampling over the CV can be achieved by standard techniques such as metadynamics or harmonic restraints [36, 21, 37]. If is the resulting bias potential, the true distribution can be obtained from the observed distribution via
Once the is sampled, the cross-sectional Marcus free energy curves are obtained by invoking the standard assumption that they are parabolic, whereby
Here and denote, respectively, the mean and variance of the conditional distribution which can be approximated from 222Interestingly, if we assume the energy gap sampling (Eq. (1)) is harmonic, equating the change in free energy between the minima of the Marcus parabola at a particular value of the CV implies that This equation relates the ET-coordinate statistics to the CV statistics, although in practice, it may not hold exactly due to asymmetry/anharmonicity in the ET, sampling insufficiency, or methodological uncertainty..
In the standard case that the sampling is performed such that is constant, using Eq. (5)-(7), we now obtain for our free energy surfaces
In the demonstration presented below, we further simplify by assuming , only because the available sampling was insufficient to resolve any -dependence from the statistical noise across runs. Nonetheless, we emphasize that the reorganization energy may, in general, be both asymmetric and depend on . (Note we assume that both the ET and CV coordinates are overdamped, with the CV relaxing on a slower timescale than the ET coordinate, allowing to be obtained from enhanced sampling [39, 19]. These assumptions will be revisited in future work).
Finally, the free energy curves obtained via the above method can be used to compute the TS and activation barrier. The intersection of the curves is given by the set
and the TS and activation barrier in the non-adiabaitc limit are given by
Adiabatic corrections to these values are discussed below.
For visualization and comparison of landscapes between systems, it is convenient to nondimensionalize both reaction coordinates. For the CV, the natural choice is
where and are determined from Eq. (4) and the tilde denotes a nondimensionalized quantity. If we plot the CV and ET coordinates on perpendicular axes (see Supplementary Material for analysis of the coordinate orthogonality), and enforce that the initial and final Marcus parabolas in each cross section lie at 0 and 1 respectively, we obtain for the ET coordinate
Note that the nondimensionalization scheme stretches the ET coordinate differently at different CV values. However, it can be easily seen that this transformation maps only the ranges of the reaction coordinates, and so barrier heights in Eq. (10b) are unaffected by the non-dimensionalization.
Once obtained, the activation barrier can be used to derive ET rate expressions. As in standard Marcus theory, we note that the thermodynamic free energy which determines the shift of the two diabatic surfaces relative to eachother, can depend on factors such as electrochemical overpotential, which only weakly effect the shape of the individual surfaces. Thus can be obtained from Eq. (10) numerically, and the non-adiabatic rate expression for a discrete transition can be determined from Fermi’s golden rule
where is a partition function and is the electronic coupling.
The rate expression for electron transfer where is a discrete state and is a continuum of states can be derived using a generalized MHC integral, integrating over the continuum
where is the density of the continuum, is the Fermi-Dirac occupancy, and is now a function not only of some global shift but also the location of the final state’s energy within the continuum.
We emphasize that the formalism presented up to this point is general, and the analytical CIET expressions of Ref. [7] can be derived upon specialization to certain electrochemical system and making the assumption that the free energy surfaces have a particular shape.
Example Application: CO2 redox. The choice of the CO2 redox system for demonstration of the above methodology is due to the relative simplicity of approximating the diabatic states. Defining the diabatic states for lithium-ion intercalation, for example, is much more difficult because the lithium ion does not change its oxidation state during the intercalation process. Instead, the electrode is reduced to accommodate the lithium ion’s charge, leading to diabatic states whose definition remains an open scientific question [12]. A simplified approach to this problem was adopted in Ref. [11] and will be extended in future work. Here, we instead demonstrate this Letter’s methodology on the supposed rate-determining step of CO2 reduction [40, 41, 42], where the diabatic states— and —are obtained via simple constraints, and where a CIET-like approach is less explored.
| -0.50 | 0.2 | 1.1 | 1.1 | 0.7 | 2.0 | 1.6 |
| 0.0 | 0.5 | 1.4 | 1.4 | 0.5 | 1.7 | 1.4 |
| 0.5 | 0.8 | 1.7 | 1.7 | 0.3 | 1.4 | 1.2 |
| 1.0 | 1.1 | 2.0 | 2.0 | 0.05 | 1.1 | 1.0 |
Electrocatalyzed CO2 reduction is of growing interest for various industrial processes and sustainability applications [43, 44], and gold and silver electrodes are noteworthy for their high product-specificity for CO [45]. Efficient ET in this system requires the presence of cations in the solution, and the mechanism of this effect has been debated [46, 47, 48]. Recent simulation results have suggested that the cations lower the barrier for CO2 adsorption, thus facilitating the ET [40].
Consistent with this picture, we take as our classical CV the -coordinate of the Carbon atom defined as the (negative) distance from the electrode surface, and take and in Eq. (1) to be the electronic states and respectively. Although other choices of CV are possible (e.g., the O–C–O bond angle; see the Supplementary Material), this -coordinate most directly captures the interfacial ion–electron coupling relevant to the CIET mechanism.
Using these definitions, we compute the diabatic free-energy surface for CO2 redox on a gold electrode in the presence of K+ ions (see Fig. 1(a) and the Supplementary Material for corresponding numerical details). Figure 2(b)-(d) shows contour and perspective views of the resulting free energy surface for several thermodynamic driving forces . From Eq. (4) we obtain and , which map to the points and respectively upon nondimensionalization. The contour plot shows relative minima in both corners corresponding to the oxidized and reduced state—the signature of a CIET-like mechanism.
It is illustrative to compare the barrier height in the 2D landscape to that which would be obtained from only considering the CV or ET coordinate alone. This comparison can be done for both the reduction and oxidation directions, and for various thermodynamic driving forces . Results of this comparison are shown in Table I. We denote by the reaction barrier that would be obtained by considering only the CV dimension, collapsing the ET dimension to the vertex of each Marcus parabolas in the manifold. We denote by the reaction barrier that would be obtained by considering only the ET dimension, taking the cross sectional Marcus parabolas at fixed CV. Finally we denote by the activation barrier obtained from the full 2D approach, which is obtained numerically from Eq. (10). Note crucially that while corresponds to a valid (although possibly higher energy) reaction path, does not correspond to a physical reaction path that satisfies the Franck Condon principle [1, 49] and is included only for illustrative purposes. We see that the CV-only approach is consistently way too small for both the reduction and oxidation directions, suggesting the importance of ET in this reaction. Though the ET-only approach is in good agreement with the full 2D approach in the reduction direction, it disagrees significantly for the oxidation direction.
Note that the evaluation of the reaction barrier at various provides a numerical method to evaluate the rate Eq. (18) above, and the application of this method to experimental reaction rates will be the subject of future work.
Adiabatic Corrections.
The landscape presented above can be modified when the reaction has adiabatic character, which is thought to be the case for CO2 redox [40, 42]. In this case, if an effective coupling were known, the adiabatic surface could be computed by diagonalizing the two-level system
Note that the true coupling value
corresponding to a nuclear coordinate R can be approximated from the Kohn-Sham cDFT orbitals according to Ref. [50], but will fluctuate throughout an AIMD trajectory. We can thus time average to approximate a mean coupling
where is the total time of the trajectory, assumed to be sufficiently large. In Eq. (17) We have first averaged along the minima in the CV direction and then employed a linear interpolation to approximate the coupling for values . In this case, as discussed in Ref. [16], the corresponding adiabatic rate expression differs only by the prefactor
Figure 2(a) shows a contour plot of the electronic coupling across the full 2D landscape, while Fig. 2(b) shows a contour plot of the computed adiabatic free energy surface adjusted according to Eq. (15). Note that the coupling values are quite large—on the order of 1–3 eV—consistent with the theoretical prediction of Ref. [16].
| 0.3 | 0.4 | 0.4 | 0.8 | 1.4 | 0.9 | ||
| 0.0 | 0.4 | 0.5 | 0.8 | 0.5 | 0.5 | 0.9 | 0.5 |
| 0.5 | 0.9 | 0.7 | 1.3 | 0.7 | 0.1 | 0.5 | 0.2 |
| 1.0 | 1.5 | 1.0 | 1.8 | 1.0 | 0.0 | 0.0 | 0.0 |
To calculate the TS, we note that in the adiabatic case, the TS is not confined to occur within the set defined in Eq. (9), and so the restriction in the minimization in Eq. (10a) must be lifted. We therefore located the TS using a minimax (lowest-saddle) search, in which all neighboring grid points were connected by edges weighted by the higher of their energies, and the smallest edge-maximum path connecting the bottom-left and top-right minima was located via a Kruskal-style union–find algorithm [51]. Furthermore, note that in contrast to the non-diabatic approximation above, the experimentally observed driving force is no longer strictly equivalent to the diabatic shift , so the latter must be adjusted analyze the former [16].
The resulting energy barriers are reported in Table II. We observe that many of the barriers are significantly lower compared to the diabatic landscape, and notably barrierless transition is obtained at an overpotential not predicted in the non-diabatic case. Furthermore, both in both the reduction and oxidation directions, there is no consistent agreement between the barriers calculated from either 1D approach and the full 2D landscape. These results provide evidence that the multidimensional, adiabatic-corrected approach presented in this Letter is necessary to accurately compute ET barrier heights in general.
In conclusion, we have introduced a general framework for constructing the 2D free energy surface of an ET reaction coupled to a classical CV in terms of constrained ab initio trajectories, which can be improved along with sampling efficiency and accurate simulation of diabatic states. Applied to an example reaction of CO2 redox, we showed that the activation barrier computed either using the ET coordinate or CV alone can be significantly misleading. Although further sampling is needed to resolve the CV-dependence of the reorganization energy, our framework readily accommodates such dependence and would be expected to yield even larger deviations from the Marcus picture. We expect that the coupled-electron transfer view presented in this Letter will provide a more accurate perspective on the activation of a wide class of electron transfer reactions in anisotropic or interfacial environments.
Acknowledgements. This work was supported by the MIT Faculty Research Innovation Fund (FRIF) through the generosity of Irv and Melinda Simon. E.A. is grateful to Ezra Alexander, Mohammadhasan Dinpajooh, Beck Hanscam, Ziyue Jiang, Matthias Kick, Abraham Nitzan, Joakin Stenlid, Leah Weisburn, Junghyun Yoon, and Chao Zhang for useful discussions.

### v3

@@
 The Marcus theory of electron transfer assumes that diabatic energy gaps are sampled from a single ensemble.
-This assumption can break down in spatially anisotropic environments, such as electrochemical interfaces or biomolecular structures, where distinct solvent ensembles arise along a collective variable describing the anisotropy.
+This assumption can break down in spatially anisotropic environments, such as Faradaic reactions at electrochemical interfaces, where distinct solvent ensembles arise along a collective variable describing the anisotropy.
 Treating this collective variable as an additional reaction coordinate linearly independent from the Marcus reaction coordinate, we develop a formalism that enables calculation of the resulting Coupled Ion–Electron Transfer (CIET) free-energy surface directly from constrained ab initio trajectories.
@@
 In CIET, this discrepancy is resolved by postulating a 2D free energy surfaces which can in general have an arbitrary structure, but in the simplest approximation is linear when cross-sectioned along the ion-transfer CV and quadratic along the Marcus ET coordinate [7].
-The result is that the CIET rate expression reduces to the Marcus–Hush–Chidsey (MHC) equation [16, 17, 18] in the limit where the reorganization energy dominates, and reduces to the BV equation in the limit where the barrier along the classical IT coordinate dominates [7].
+The result is that the CIET rate expression reduces to the Marcus–Hush–Chidsey (MHC) equation [16, 17, 18] with an added IT prefactor in the limit where the reorganization energy dominates, and reduces to the BV equation in the limit where the barrier along the classical IT coordinate dominates [7].
 The fact that viewing an electron transfer reaction as coupled to an anisotropic CV not only explains experimental reaction rates but also unifies physical equations correct in different regimes suggests the utility of such an approach, and suggests the possibility that a related framework can be generalized to symmetry-broken CVs other than IT at intercalation cathodes, finding broader applications across condensed phase physical chemistry [19, 20].
 Practical analyses using these multidimensional surfaces may also become more feasible with concurrent improvements in sampling methods and machine-learned potentials [21, 22, 23, 24, 25, 26].
-We note that ab initio computation of CIET parameters has recently been pioneered by Stenlid et al.
-[11] on a model lithium-cobalt-oxide battery interface.
-While remarkably effective in qualitatively reproducing experimental kinetic data, these calculations were conducted at 0 K, leaving entropic contributions to the free energy unaddressed.
-Furthermore, the study stopped short of computing the full 2D landscape, making assumptions, for example, that the reorganization energy was independent of the IT coordinate, and that adiabatic corrections to the landscape were insignificant.
-Here we present a formalism to compute finite-temperature 2D free-energy surfaces for ET coupled to a classical CV, expressed entirely in terms of quantities derived from constrained ab initio trajectories.
+We note that ab initio computation of CIET parameters has recently been explored by Stenlid et al.
+on a model lithium-cobalt-oxide battery interface [11].
+While their reported parameters reproduced trends observed in experimental current-voltage relationships, we noticed several methodological inconsistencies in their study (see Supplementary Material for details).
+Here we present a physically consistent formalism to compute finite-temperature 2D free-energy surfaces for ET coupled to a classical CV, expressed entirely in terms of quantities derived from constrained ab initio trajectories.
 We incorporate adiabatic corrections to the landscape as well, which prove to be significant [16].
@@
 Consider a simulation of atoms with the electronic state fixed to , where denotes the nuclear configuration.
-Let denote the nuclear trajectory under this constraint.
+Let denote a nuclear trajectory under this constraint.
 At steady state, the probability of finding the system in configuration is given by the time-independent distribution .
@@
 E.A.
-is grateful to Ezra Alexander, Mohammadhasan Dinpajooh, Beck Hanscam, Ziyue Jiang, Matthias Kick, Abraham Nitzan, Joakin Stenlid, Leah Weisburn, Junghyun Yoon, and Chao Zhang for useful discussions.
+is grateful to Ezra Alexander, Mohammadhasan Dinpajooh, Beck Hanscam, Ziyue Jiang, Matthias Kick, Abraham Nitzan, Joakim Stenlid, Leah Weisburn, Junghyun Yoon, and Chao Zhang for useful discussions.